OpenStudy (anonymous):

Find the arc length of the curve x= (1/3)t^3 y= (1/2)t^2 from zero is less and equal to t and t is greater than or equal to 1

OpenStudy (anonymous):

\[x=\frac{ 1 }{ 3 }t ^{3}\] \[y=\frac{ 1 }{ 2 }t ^{2}\]

OpenStudy (ybarrap):

The Formula for Arclength is: $$ \large s=\int _{{a}}^{{b}}{\sqrt {1+[f'(x)]^{2}}}\,dx. $$ Have you used it?

OpenStudy (anonymous):

but now i am doing parametric

OpenStudy (ybarrap):

Ok, so you'll need this one. Have you used it? $$ \large s=\int _{{a}}^{{b}}{\sqrt {[X'(t)]^{2}+[Y'(t)]^{2}}}\,dt. $$

OpenStudy (ybarrap):

1st determine $$ \large { \cfrac{dx(t)}{dt}\text{&} \cfrac{dy(t)}{dt}\\} $$

OpenStudy (anonymous):

i did dx/dt= t^2 dy/dt= t

OpenStudy (anonymous):

what confused me is the integral part i keep getting